OpenStudy (anonymous):

Show how you solve the following two equations: x + 3 = 18

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (leilanilane):

X + 3 = 18 Is the same thing as 18 - 3 = x So x = 15

OpenStudy (ipwnbunnies):

Yes, since 3 is being added to x, you'll want to subtract it to the other side.

OpenStudy (ipwnbunnies):

This is in English, btw.

OpenStudy (leilanilane):

As simple as it gets lol

OpenStudy (anonymous):

I don't care if its in English buns:)

OpenStudy (ipwnbunnies):

Ok lol. Where's the second equation?

OpenStudy (leilanilane):

Does no one know that answer to my question :/ Lol

OpenStudy (ipwnbunnies):

Study groups? I don't think you can do that. Unless you make a question, then close it right when you make it.

OpenStudy (anonymous):

x + 3/4=2 1/4

OpenStudy (leilanilane):

No i read on the openstudy blog that they made a feature where u can add study groups and topics so unless they took that down...

OpenStudy (ipwnbunnies):

Oh, idk.

OpenStudy (anonymous):

???

OpenStudy (ipwnbunnies):

For this question, we want to rewrite these fractions as fractions with the same denominator. We also want to get rid of the mixed number on the right side.

OpenStudy (ipwnbunnies):

\[2 \frac{1}{4} = \frac{2*4+1}{4}\]

OpenStudy (anonymous):

I aint got no calcumaltor

OpenStudy (anonymous):

;(

OpenStudy (ipwnbunnies):

calculator* You don't know 2*4+1? What's 2*4

OpenStudy (anonymous):

calcumalator* 8

OpenStudy (ipwnbunnies):

Lol && Now add 1. :)

OpenStudy (anonymous):

9!

OpenStudy (anonymous):

right =.=

OpenStudy (ipwnbunnies):

\[2 \frac{1}{4} = \frac{9}{4}\] \[x + \frac{3}{4} = \frac{9}{4}\]

OpenStudy (ipwnbunnies):

We changed the right side of the equation. Now, when adding and subtracting fractions, they must have the same denominator, or bottom number. Then, we do the math on the top numbers.

OpenStudy (ipwnbunnies):

Since both fractions have '4' in the bottom number, we can solve this easily. X is being added to 3/4. So, what do we do to get x by itself?

OpenStudy (anonymous):

I don't know!:(

OpenStudy (anonymous):

hey don't look at me I confused

OpenStudy (anonymous):

cornfusededed*

OpenStudy (ipwnbunnies):

Since x is being added to 3/4, you need to subtract. It's the opposite of addition. We subtract 3/4 from both sides to get x by itself.

OpenStudy (anonymous):

oh. im less cornfusedededed.

OpenStudy (ipwnbunnies):

Good. \[x = \frac{9}{4} - \frac{3}{4}\] As I said, when adding or subtracting fractions. When they have similar bottom numbers, you can add and subtract the top numbers. We don't add and subtract the bottom numbers. So, what would the right side be?

OpenStudy (anonymous):

oh crap I don't know im stupid and blonde. I think I can talk to my brother about it if you want to go to bed:)

OpenStudy (ipwnbunnies):

No, we're almost there. What's 9-3? We do the math on the top numbers when the bottoms are the same.

OpenStudy (anonymous):

ok...

OpenStudy (ipwnbunnies):

\[x = \frac{9}{4} - \frac{3}{4}\] \[x = \frac{6}{4} = \frac{3}{2}\]

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

thanx